OpenStudy (anonymous):

Order the numbers from least to greatest. 5^-3, 4/7 , √5 , 3/8 , 0.003

OpenStudy (anonymous):

this is my answer: 5^-3 , 0.003 , √5 , 4/7 , 3/8 am i right? @RoseDryer @kausarsalley

OpenStudy (anonymous):

can you change the other numbers to decimal??

OpenStudy (anonymous):

4/7 = 0.57142857142857. 3/8 = 0.375. √5= 2.5. 0.003. and i'm not sure how to make 5^-3 a decimal

OpenStudy (anonymous):

@kausarsalley

OpenStudy (anonymous):

are you aware \[5^{-3} = \frac{ 1 }{ 5^3 }\]

OpenStudy (anonymous):

does that help to change it into a decimal??

OpenStudy (anonymous):

0.008 @kausarsalley

OpenStudy (anonymous):

yes...so that is it..

OpenStudy (anonymous):

can you now rearrange it??

OpenStudy (anonymous):

0.008, 0.003, 0.57142857142857, 0.375, 2.5 is that right???

OpenStudy (anonymous):

@kausarsalley

OpenStudy (anonymous):

you have mixed them up..... 0.003 is smaller than 0.008 and also 0.375 is smaller than 0.571.....

OpenStudy (anonymous):

so it's 0.003, 0.008, 0.375, 0.57142857142857, 2.5 ?

OpenStudy (anonymous):

@kausarsalley

OpenStudy (anonymous):

yes.......that is correct!......(:

OpenStudy (anonymous):

Thank you so much!! I seriously love you!! @kausarsalley

OpenStudy (anonymous):

you are welcome......(: